# Category Adaptation Meets Projected Distillation in Generalized Continual Category Discovery

### v2

@@ Abstract
Generalized Continual Category Discovery (GCCD) tackles learning from sequentially arriving, partially labeled datasets while uncovering new categories. Traditional methods depend on feature distillation to prevent forgetting the old knowledge. However, this strategy restricts the model’s ability to adapt and effectively distinguish new categories. To address this, we introduce a novel technique integrating a learnable projector with feature distillation, thus enhancing model adaptability without sacrificing past knowledge. The resulting distribution shift of the previously learned categories is mitigated with the auxiliary category adaptation network. We demonstrate that while each component offers modest benefits individually, their combination – dubbed CAMP (Category Adaptation Meets Projected distillation) – significantly improves the balance between learning new information and retaining old. CAMP exhibits superior performance across several GCCD scenarios with or without exemplars. Additionally, CAMP translates to a well-established Class Incremental Learning setting, achieving state-of-the-art results.
@@ 1 Introduction
To address this issue, we represent categories as centroids in the latent space of the feature extractor and train an auxiliary network to predict the drift of centroids, which we call category adaptation. This enables the model to recover the actual positions of old clusters and improve the overall accuracy of the Nearest Centroid Classifier (NCC). Similarly, works [17, 41] analyze the semantic drift of old categories in closed-world, Continual Learning scenarios. Here, our key observation is that, while centroid adaptation demonstrates overall solid performance, its true effectiveness is realized only through combining it with feature distillation using a projection network, as depicted in Fig. 1. That allows the technique to excel. Building upon these findings, we propose to leverage this combined effect for GCCD by introducing a method called CAMP – Category Adaptation Meets Projected distillation. CAMP presents state-of-the-art performance across GCCD scenarios and exemplar-free Class Incremental Learning. To summarize, the main contributions of our work are as follows:
  • We study the interplay between knowledge distillation and category adaptation for the problem of Generalized Continual Category Discovery (GCCD). We demonstrate that they address distinct or even opposing challenges when considered separately. However, when combined, they complement each other significantly, reduce forgetting, and improve overall results.
  • Based on those insights, we propose a novel method for generalized continual category discovery dubbed CAMP. Built on GCD [35], CAMP improves it for continual learning by adapting categories between tasks and performing knowledge distillation with two separate projectors.
  • We evaluate our method on numerous GCCD and Class Incremental Learning (CIL) datasets. With a series of experiments, we show that CAMP achieves state-of-the-art results with and without exemplars in different scenarios.
@@ 2 Related work
Continual classes discovery combines the NCL and CL fields. In [32], learning of the novel classes is framed incrementally: that is, in the first stage, the model learns from supervised data, and in the following stages, novel classes are learned without any labels. A similar setting was explored in [18] with the solution that uses pseudo-latent supervision, feature distillation, and a mutual information-based regularizer. However, their setting assumes no class overlap between the initial and incremental training stages, which was later relaxed by various to which we refer as Generalized Continual Category Discovery - GCCD [44, 20, 38, 43]. Grow and Merge (GM)[43] keeps two models: a static one that maintains knowledge of old classes and a dynamic one trained using self-supervision to discover novel classes. Both models are then merged into one when the new task arrives. In IGCD [44] a non-parametric classifier and a small buffer of exemplars from previous tasks to decrease forgetting. Another approach, PA [20] also uses exemplars and combines them with the proxy anchors [21] known from deep metric learning. Contrary to previous approaches [43, 44, 20] we focus on the most challenging exemplar-free setting. MetaGCD [38] needs to store the data from the first task, which we also can avoid, by proposing a purely continual model. We show that the proposed method improves in both settings with and without exemplars over existing approaches.
@@ 4.1 Experimental setup
We evaluate all methods on five commonly used datasets in CL and GCD: CIFAR100 [24] consisting of 60,000 images of 100 classes, Stanford Cars [23] consisting of 16,185 images of cars, CUB200 [36] representing 11,788 images of bird species, FGVCAircraft [26] dataset consisting 10,200 images of planes and DomainNet [28] representing data from 6 domains. We split datasets into five equal tasks, with the exception of Stanford Cars (4 tasks) and DomainNet (6 tasks). We set a different domain for each task of DomainNet, which allows us to simulate scenarios with domain shifts. Following [35, 37], for each dataset, we set the ratio of known to novel classes as 4:1. For known classes, we set the ratio of labeled data samples to unlabeled data as 1:1. We report results from the test set on known, novel, and as well as all classes together. Note that the number of tasks $N$, percentage of novel classes, and percentage of labeled samples for known classes are hyperparameters of the GCCD setting.
@@
As a feature extractor $\mathcal{F}$ for all methods we use ViT-small [8] model pretrained with DINO [4] on ImageNet [7]. It has 12 transformer blocks, 12 heads per block, and an embedding size of 384. Following GCD [35] we freeze the first 11 blocks for all methods. For a fair comparison, we train all methods for 100 epochs in each task using AdamW optimizer with a starting learning rate of $0.1^{3}$. The exception is PA method, for which we set it to $0.1^{4}$ but increase the learning rate for proxies 100 times following the original work. For all methods, we utilize warm-up and cosine annealing scheduling, which decays the learning rate to $0.1^{6}$. As the distiller $\phi$ we utilize a 2-layer MLP network consisting of 384 neurons in each layer with batch normalization followed by ReLU activation function after the first layer. As the adapter $\psi$, we utilize a linear layer consisting of 384 neurons. For CAMP without exemplars, we set $\alpha$ to 0.5; for CAMP with exemplars, we use $\alpha=0.1$.
@@
Selection of baselines For the simplest baseline, we fine-tune GCD [35] method. However, as it was not designed for continual learning use, we couple it with a popular CL regularization-based methods: (1) weight regularization method - EWC [22] (GCD+EWC) where we add a linear head and replace SupCon loss with cross-entropy; (2) feature distillation method [25, 13] (GCD+FD). In GCD+FD we distill features $\mathcal{F}(x)$. Moreover, we combine GCD with experience replay [30] (GCD+ER), where for incremental tasks, we add a buffer that stores random images and labels of past known classes. For all the above methods, we perform semi-supervised k-means [35] algorithm to find centroids and utilize the NCC classifier to classify input images. Additionally, we use SimGCD [37], which improves GCD by utilizing a learnable classifier and adding additional cross-entropy-based losses for training it. We also test MetaGCD[38] that utilizes meta-learning and PA [20] method, which trains the feature extractor using supervised proxy anchor loss, maintains a buffer of exemplars, and uses feature distillation to alleviate forgetting. Lastly, we evaluate IGCD [44] that utilizes a density-based support set selection and replay buffer to mitigate forgetting.
We use original implementations of all methods except PA where the code was not provided, so we re-implemented it following the original work. We set hyperparameters of all baseline methods following the original works. We describe all methods in detail in the Supplementary Material C.
@@ 4.2 Comparison to baselines
To better examine how all accuracy changes in incremental steps, we plot averages of all accuracy after each task in Fig. 4. CAMP achieves the highest accuracy from the second task, and the results are consistent with and without exemplars. Next, we analyze the plasticity and forgetting of methods in Fig. 5 to explain the high performance of CAMP. GCD is the most plastic method. However, it has the highest forgetting as it has no form of regularization. On the other hand, IGCD is the most stable as it achieves the lowest forgetting. CAMP has good plasticity and forgetting, allowing it to achieve state-of-the-art results. We can also observe that exemplars prevent methods like GCD, IGCD, PA, and CAMP from forgetting. That is intuitive as the replay buffer allows them to recall the past categories’ data. Interestingly, each method forms its own cluster in presented plots, informing that they have different but specific characteristics for continual learning scenarios.
Figure 4: Average accuracy of methods after each task. CAMP achieves better results than competitors on CIFAR100 during the continual learning session of five tasks, both with and without exemplars.
Figure 5: Compared to other methods, CAMP achieves better forgetting and plasticity tradeoffs, resulting in good overall average accuracy. The bottom right is the sweet spot.
4.3 CAMP when all data is labeled - Class Incremental Learning
Class incremental learning [27] is a special case of continual Generalized Category Discovery, where each data sample is labeled, and all classes are known. We test CAMP against exemplar-free state-of-the-art methods and train the feature extractor (ResNet-18) from scratch. We train methods for $N$ equal tasks on CIFAR100, Domainnet, and Imagenet-Subset. For training CAMP we remove the self-supervised component from the loss function as all data is labeled. For baselines, we utilize popular exemplar-free CIL methods using their implementations in well-established benchmarks (FACIL [27], PyCIL [48]) and set their hyperparameters to default. We implement CAMP in FACIL and will publish the code upon the acceptance of the manuscript. For details, please refer to Supplementary Material B. We report average incremental accuracy [27], which is the average of task agnostic accuracies measured after each task on tasks seen so far. We provide results in Tab. 2.
@@ 4.4 Method Analysis
Estimating number of categories in a task. In order to estimate the number of novel categories in each task, we run an elbow method in every task. We present results in Fig. 7. It is clear that accuracy on the labeled dataset peaks when the number of clusters for k-means algorithm is equal to the ground truth number of categories.
@@
CAMP vs. GCD under different setting conditions. We examine the impact of setting parameters (number of novel classes and fraction of labeled data) on results achieved by CAMP and GCD. We set novel to known ratios as 2:8, 4:6, 6:4, 8:2 and fraction of labeled data as 0.25, 0.5, 0.75, 1.0. We provide final all accuracy and average forgetting on StanfordCars in Fig. 11. We can observe that accuracy decreases when increasing the number of novel classes or when decreasing the amount of labeled data. Forgetting behaves similarly, which is understandable - with lower accuracy, there is less to forget. However, in each case, CAMP outperforms GCD in every scenario, proving that our combination of distillation and category adaptation is robust to different scenarios.
@@ 5 Summary
In this work, we study an interplay between knowledge distillation techniques and centroid adaptation in the problem of Generalized Continual Category Discovery and CIL. We find out that combining knowledge distillation through a neural projector followed by centroid adaptation is crucial for achieving high plasticity and low forgetting. Based on this finding, we propose a new method, dubbed CAMP. It utilizes projected knowledge distillation to regularize the network and prevent forgetting while training from the stream of non-i.i.d data. Moreover, CAMP utilizes a centroid adaptation method to estimate better actual positions of centroids representing past categories. CAMP uses the nearest centroid classifier to classify data samples.
In the series of experiments, we compared CAMP to existing state-of-the-art Generalized Continual Category Discovery methods and Class Incremental Learning methods on several well-established datasets. Our approach consistently outperforms these strong baselines in terms of task-agnostic accuracy while achieving better plasticity/forgetting trade-offs.

### v4

@@ Abstract
Generalized Continual Category Discovery (GCCD) tackles learning from sequentially arriving, partially labeled datasets while uncovering new categories. Traditional methods depend on feature distillation to prevent forgetting the old knowledge. However, this strategy restricts the model’s ability to adapt and effectively distinguish new categories. To address this, we introduce a novel technique integrating a learnable projector with feature distillation, thus enhancing model adaptability without sacrificing past knowledge. The resulting distribution shift of the previously learned categories is mitigated with the auxiliary category adaptation network. We demonstrate that while each component offers modest benefits individually, their combination – dubbed CAMP (Category Adaptation Meets Projected distillation) – significantly improves the balance between learning new information and retaining old. CAMP exhibits superior performance across several GCCD and Class Incremental Learning scenarios. The code is available on Github.
@@ 1 Introduction
To address this issue, we represent categories as centroids in the latent space of the feature extractor and train an auxiliary network to predict the drift of centroids, which we call category adaptation. This enables the model to recover the actual positions of old clusters and improve the overall accuracy of the Nearest Centroid Classifier (NCC). Similarly, works [21, 50] analyze the semantic drift of old categories in closed-world, Continual Learning scenarios. Here, our key observation is that, while centroid adaptation demonstrates overall solid performance, its true effectiveness is realized only through combining it with feature distillation using a projection network, as depicted in Fig. 1. That allows the technique to excel. Building upon these findings, we propose to leverage this combined effect for GCCD by introducing a method called CAMP – Category Adaptation Meets Projected distillation. CAMP presents state-of-the-art performance across GCCD scenarios and exemplar-free Class Incremental Learning (CIL).
To summarize, the main contributions of our work are as follows. 1: We study the interplay between knowledge distillation and category adaptation for the problem of Generalized Continual Category Discovery (GCCD). We demonstrate that they address distinct or even opposing challenges when considered separately. However, when combined, they complement each other significantly, reduce forgetting, and improve overall results. 2: Based on those insights, we propose a novel method for generalized continual category discovery dubbed CAMP. Built on GCD [43], CAMP improves it for continual learning by adapting categories between tasks and performing knowledge distillation with two separate projectors. 3: We evaluate our method on numerous GCCD and Class Incremental Learning (CIL) datasets. With a series of experiments, we show that CAMP achieves state-of-the-art results with and without exemplars in different scenarios.
@@ 2 Related work
Continual classes discovery combines the NCL and CL fields. In [39], learning of the novel classes is framed incrementally: that is, in the first stage, the model learns from supervised data, and in the following stages, novel classes are learned without any labels. A similar setting was explored in [22] with the solution that uses pseudo-latent supervision, feature distillation, and a mutual information-based regularizer. However, their setting assumes no class overlap between the initial and incremental training stages, which was later relaxed by various to which we refer as Generalized Continual Category Discovery - GCCD [53, 25, 47, 52]. Grow and Merge (GM)[52] keeps two models: a static one that maintains knowledge of old classes and a dynamic one trained using self-supervision to discover novel classes. Both models are then merged into one when the new task arrives. In IGCD [53] a non-parametric classifier and a small buffer of exemplars from previous tasks to decrease forgetting. PA [25] combines exemplars and proxy anchors [26] known from deep metric learning. Contrary to previous approaches [52, 53, 25] we focus on the most challenging exemplar-free setting. MetaGCD [47] needs to store the data from the first task, which we avoid, by proposing a purely continual model.
@@ 4.1 Experimental setup
We evaluate all methods on five datasets: CIFAR100 [29], Stanford Cars [28], CUB200 [44], FGVCAircraft [32] and DomainNet [34]. We split datasets into five equal tasks, with the exception of Stanford Cars (4 tasks) and DomainNet (6 tasks). Following [40] we set a different domain for each task of DomainNet, to simulate domain shifts. Following [43, 46], for each dataset, we set the ratio of known to novel classes as 4:1. For known classes, we set the ratio of labeled data samples to unlabeled data as 1:1. As the feature extractor $\mathcal{F}$ for all methods we use ViT-small [11] model pretrained with DINO [7] on ImageNet [10]. Following GCD [43] we freeze the first 11 blocks for all methods. We train all methods for 100 epochs in each task using AdamW optimizer with a starting learning rate of $0.1^{3}$ and cosine annealing scheduling. The exception is PA method, for which we set it to $0.1^{4}$ but increase the learning rate for proxies 100 times following the original work. As the distiller $\phi$ we utilize a 2-layer MLP network consisting of 384 neurons and ReLU activation. As the adapter $\psi$, we utilize a linear layer consisting of 384 neurons. For CAMP without exemplars, we set $\alpha$ to 0.5; for CAMP with exemplars, we use $\alpha=0.1$.
@@
Selection of baselines For the simplest baseline, we fine-tune GCD [43] method and couple it with popular CL regularization-based methods: (1) weight regularization method - EWC [27] (GCD+EWC); (2) feature distillation method [30, 17] (GCD+FD). In GCD+FD we distill features $\mathcal{F}(x)$. Moreover, we combine GCD with experience replay [37] (GCD+ER), where for incremental tasks, we add a buffer that stores random images of past known classes. For all the above methods, we perform semi-supervised k-means [43] algorithm to find centroids and utilize the NCC classifier for classification. Additionally, we use SimGCD [46], which improves GCD by utilizing a learnable classifier and adding additional cross-entropy-based losses for training it. We also test MetaGCD[47] that utilizes meta-learning and PA [25] method, which trains the feature extractor using supervised proxy anchor loss, maintains a buffer of exemplars, and uses feature distillation to alleviate forgetting. Lastly, we evaluate IGCD [53] that utilizes a density-based support set selection and replay buffer to mitigate forgetting. We describe all methods in detail in the Supplementary Material C.
@@ 4.2 Comparison to baselines
To better examine how the accuracy changes in incremental steps, we plot averages of all accuracy after each task in Fig. 4 (left). CAMP achieves the highest accuracy from the second task, and the results are consistent with and without exemplars. Next, we analyze the plasticity and forgetting of all methods in Fig. 4 (right). GCD is the most plastic method. However, it has the highest forgetting as it has no form of regularization. On the other hand, IGCD is the most stable as it achieves the lowest forgetting. CAMP has good plasticity and forgetting, allowing it to achieve state-of-the-art results. We can also observe that exemplars prevent methods like GCD, IGCD, PA, and CAMP from forgetting. That is intuitive as the replay buffer allows them to recall the past categories’ data. Interestingly, each method forms its own cluster in presented plots, informing that they have different but specific characteristics for continual learning scenarios.
Figure 4: Results of GCCD methods on CIFAR100 during the continual learning session of five tasks. We report average accuracy after each task (left); plasticity and forgetting (right). CAMP achieves better accuracy and than competitors.
Class Incremental Learning results Class incremental learning [33] is a special case of GCCD, where each data sample is labeled, and all classes are known. We test CAMP against exemplar-free state-of-the-art methods and train the feature extractor (ResNet-18) from scratch. We train methods for $N$ equal tasks on CIFAR100, Domainnet, and Imagenet-Subset. For training CAMP we remove the self-supervised component from the loss function as all data is labeled. For baselines, we utilize popular exemplar-free CIL methods using their implementations in well-established benchmarks (FACIL [33], PyCIL [57]) and set their hyperparameters to default. We implement CAMP in FACIL and will publish the code upon the acceptance of the manuscript. For details, please refer to Supplementary Material B. We report average incremental accuracy [33], which is the average of task agnostic accuracies measured after each task on tasks seen so far. We provide results in Tab. 2.
@@ 4.3 Method Analysis
Ablation study We perform ablation study on CUB200 and FGVCAircraft datasets. We present results in Tab. 3. We test CAMP without: $\mathcal{L}_{SL}$, $\mathcal{L}_{SSL}$, knowledge distillation through the projector loss ($\mathcal{L}_{KD}$) and without centroid adaptation. We can see that having all of these elements is crucial for obtaining good accuracy on known and novel classes (62.6% and 44.2% respectively on CUB200). Interestingly, without the $\mathcal{L}_{SSL}$ CAMP achieves 2.0% points better accuracy for known classes but 2.8% points lower accuracy for novel classes.
Table 3: Ablation study of CAMP.
CUB200 FGVCAircraft
   $\mathcal{L}_{SL}$ $\mathcal{L}_{SSL}$ $\mathcal{L}_{KD}$ Cent. adapt. Known Novel Known Novel
   ✗ $\checkmark$ $\checkmark$ $\checkmark$ 29.9$\pm$1.6 12.1$\pm$1.1 19.0$\pm$0.9 18.8$\pm$1.9
   $\checkmark$ $\checkmark$ $\checkmark$ 64.6$\pm$2.3 41.4$\pm$3.5 43.5$\pm$1.8 16.6$\pm$2.0
   $\checkmark$ $\checkmark$ $\checkmark$ 30.9$\pm$2.6 32.9$\pm$2.6 24.7$\pm$1.2 19.8$\pm$2.2
   $\checkmark$ $\checkmark$ $\checkmark$ 36.5$\pm$1.9 28.7$\pm$1.7 29.6$\pm$1.7 24.2$\pm$2.8
   $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 62.6$\pm$2.4 44.2$\pm$3.7 42.0$\pm$1.9 31.5$\pm$3.5
Estimating number of categories in a task. We estimate the number of novel categories in each task using the elbow method. We present results in Fig. 6. It is clear that accuracy on the labeled dataset peaks when the number of clusters for k-means algorithm is equal to the ground truth number of categories.
@@
CAMP vs. GCD under different setting conditions. We examine the impact of setting parameters (number of novel classes and fraction of labeled data) on results achieved by CAMP and GCD. We set novel to known ratios as 2:8, 4:6, 6:4, 8:2 and fraction of labeled data as 0.25, 0.5, 0.75, 1.0. We provide final all accuracy and average forgetting on StanfordCars in Fig. 10. We can observe that accuracy decreases when increasing the number of novel classes or when decreasing the amount of labeled data. Forgetting behaves similarly, which is understandable - with lower accuracy, there is less to forget. However, in each case, CAMP outperforms GCD, proving that our combination of distillation and category adaptation is robust to different scenarios.
@@ 5 Summary
In this work, we study an interplay between knowledge distillation techniques and centroid adaptation in the problem of GCCD and CIL. We find out that combining knowledge distillation through a neural projector followed by centroid adaptation is crucial for achieving high plasticity and low forgetting. Based on this finding, we propose a new method, dubbed CAMP. It utilizes projected knowledge distillation to regularize the network and prevent forgetting. Moreover, CAMP utilizes a centroid adaptation method to estimate better actual positions of centroids representing past categories. CAMP uses the nearest centroid classifier to classify data samples achieving state-of-the-art results in different scenarios.